OpenStudy (anonymous):

If cosx cos(pi/7) + sinx sin(pi/7) = -(sqrt(2))/2, then x can equal...? (Check all that apply) (Will give medal & fan!) A. pi/4 + pi/7 + 2npi B. 5pi/4 + pi/7 + 2npi C. 7pi/4 + pi/7 + 2npi D. 3pi/4 + pi/7 + 2npi

OpenStudy (anonymous):

A straight up answer would be much appreciated, but an explanation is okay as well

OpenStudy (anonymous):

@whpalmer4 when you get the chance can you help me?

OpenStudy (whpalmer4):

I think the trig identity that will help here is this: \[\cos(u-v) = \cos u \cos v + \sin u \sin v\]

OpenStudy (anonymous):

How will that help me? I don't understand it. Like I see how the equation matches up with it but I just don't understand.

OpenStudy (whpalmer4):

well, let's look at the first one. if \[x = \frac{\pi}{4} + \frac{\pi}7\]and we assume that the matching up means that \[\cos(x-\frac{\pi}{7}) = \cos x\cos\frac{\pi}7 +\sin x\sin\frac{\pi}7 = -\frac{\sqrt{2}}2\] that means that \(\cos(x-\frac{\pi}7) = \cos(\frac{\pi}{4}+\frac{\pi}{7}-\frac{\pi}{7}) =\cos(\frac{\pi}{4})= -\frac{\sqrt{2}}{2}\)

OpenStudy (whpalmer4):

but does \(\cos(\frac{\pi}{4}) = -\frac{\sqrt{2}}{2}\)?

OpenStudy (anonymous):

I don't think so, no.

OpenStudy (whpalmer4):

okay, so A is not a valid choice. that pi/7 stuff looks scary until you realize that it disappears in the wash, so to speak...

OpenStudy (anonymous):

So would it be 3pi/4 and 5pi/4?

OpenStudy (whpalmer4):

Yes, that appears to be correct. Good job!

OpenStudy (anonymous):

Yay! Thank you!